[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jul 31, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Aug 1 20:29:59 UTC 2017


Here is today's summary of NHC computer operations:
--- PS/MO

----------

TAFB:
1.) NWPS EPAC last several days consistently lower than WW3 by several ft.  
Appear to be a problem with swell information not coming through the  
boundaries. Doug Gaer, Chris Mellow, Craig Mattock and Pablo to look into  
it.
HSU:
1.) Wind speed probabilities did not update on website with the special 12z  
adv this morning. Matt O looking into it.

2.) Over the weekend HSU has problems saving .spf files. It appears to be a  
permissions issue related to the NetApp migration. Matt Sardi working on it  
along with Doug Gaer.

3.) Similar issues when creating outlook graphic vgf diles. Matt and Doug  
fixed when it happen but looking further into this problem. Also related to  
NetApp migration.

4.) GOES16 meso sectors reportedly coming in late in NAWIPS. Monica to look  
into this.

Other items:
1.) Major NetApp migration took place Friday. Some lingering issues with  
data gaps and permissions over the weekend related to this. But Matt and  
Doug are working on it.
NHC RFCs:
1.) RFC for NetApp migration completed. But Doug and Matt dealing with  
lingering issues.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170801/49161646/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list